# Diffusive and hydrodynamic magnetotransport around a density perturbation in a two-dimensional electron gas

### v1

@@ III.4 Analytical approximations
Being the fourth order, the viscous hydrodynamics Eq. (34) is much more complicated than the diffusive transport one (i.e., the same equation with $l_{G}=0$). We have not succeeded in finding its full analytical solutions for either of the two studied density profiles, Eqs. (12) and (13). An approximate solution can be derived if the three relevant length scales are well separated:
@@
In Fig. 5 we compare these analytical predictions with numerical results (the solid line, see the figure caption). The quantity plotted is the dimensionless combination $r\delta\psi(r)\equiv r[r+\psi(r)]/R_{\mathrm{no}}^{2}$ whose asymptotic value is $\tilde{\lambda}/R_{\mathrm{no}}^{2}$. The numerical results behave as expected, approaching the dashed line [Eq. (43)] at large $r$ and the dotted line [Eq. (90)] at small $r$. All the lines converge together at $R_{\mathrm{no}}\lesssim r\lesssim l_{G}$, validating our matching method. Given the modestly large parameter $l_{G}/R_{\mathrm{no}}=10$ used in this calculation, the agreement can be considered very good, especially for the real part of the stream function, Fig. 5(a). To improve the agreement for the imaginary part, we chose the second term in Eq. (58) to be $3iR_{\mathrm{no}}^{2}$, which is $12\%$ of the total $\Im\mathrm{m}\,\tilde{\lambda}$.
In Fig. 6 we show the 2D plots of the stream function at intermediate distances $a\ll r\lesssim l_{G}$. The zoomed-in view of the no-go region [Fig. 6(b)] reveals spiral-like pattern of the current flow, similar to the diffusive case, Fig. 3.
Figure 6: False color and contour plots of the stream function in the viscous regime for $\beta=3$, $l_{G}/R_{\mathrm{no}}=10$. (a) $\Psi(x,y)$ computed using Eq. (90). (b) Enlarged view of the no-go region, using Eq. (92).
@@ IV Discussion
The exponentially small current inside the no-go region exhibits an intriguing spiral pattern. The spiraling is driven by the Lorentz force gradient, so it has a different origin from vortical flow that can arise in viscous hydrodynamics [68, 69, 41, 70, 52] even in the absence of the magnetic field. The twisting of the current lines at short distances has an effect on the asymptotic behavior. It changes the angular orientation of the Landauer dipole, i.e., the correction to the uniform current at large distances.
The Landauer dipole and the effective obstacle size we have investigated are the simplest and the most robust characteristics. They can be extracted already from STP maps collected at a moderate resolution. By varying $B$ and tracking how the no-go radius grows, one can in principle distinguish diffusive and hydrodynamic behavior and estimate the electron viscosity. High-resolution scans [21] have been shown to also resolve a small anomaly of the electrochemical potential predicted to occur [71, 72] at the distance of one cyclotron diameter $2R_{c}$ from the hard-wall barrier. In this regard, our classical continuum medium theory has a number of limitations. Both the diffusive and the hydrodynamic approximations become inaccurate at short distances. They do not account for features on the scale of the cyclotron radius where the motion is ballistic. These approaches also fail at low temperatures and long distances, where quantum localization — the hallmark of the QHE — becomes important. To remedy these shortcomings, self-consistent solutions of the Boltzmann kinetic equation and quantum transport simulations in a realistic potential profile may be necessary. The scaling rules we have derived may provide a useful reference for these more ambitious calculations.
@@ Acknowledgements
We are grateful to Z. Krebs and V. Brar for prior collaboration on Ref. [21] that inspired this study.
@@ Appendix A Exact solutions for the diffusive transport
$\lambda=-\frac{\Gamma\left(1-\frac{2}{\beta}\right)}{\Gamma\left(1+\frac{2}{\beta}\right)}\,\frac{\Gamma\left(-\frac{i}{\alpha\beta}+\frac{1}{\beta}\right)}{\Gamma\left(-\frac{i}{\alpha\beta}-\frac{1}{\beta}\right)}\,a^{2}.$ (73)

### v2

@@ III.4 Analytical approximations
Being the fourth order, the viscous hydrodynamics Eq. (38) is much more complicated than the diffusive transport one (the same equation with $l_{G}=0$). We have not succeeded in finding its full analytical solution for either of the two studied density profiles, Eqs. (12) and (13). An approximate solution can be derived if the three relevant length scales are well separated:
@@
In Fig. 5 we compare these analytical predictions with numerical results (the solid line, see the figure caption). The quantity plotted is the dimensionless combination $r\delta\psi(r)\equiv r[r+\psi(r)]/R_{\mathrm{no}}^{2}$ whose asymptotic value is $\tilde{\lambda}/R_{\mathrm{no}}^{2}$. The numerical results behave as expected, approaching the dashed line [Eq. (47)] at large $r$ and the dotted line [Eq. (97)] at small $r$. All the lines merge together at $R_{\mathrm{no}}\lesssim r\lesssim l_{G}$, validating our matching method. Given that the parameter $l_{G}/R_{\mathrm{no}}=10$ used in this calculation is only modestly large, the agreement can be considered good, especially for the real part of the stream function, Fig. 5(a). To improve the agreement for the imaginary part, we chose the second term in Eq. (62) to be $3iR_{\mathrm{no}}^{2}$, which is $12\%$ of the total $\Im\mathrm{m}\,\tilde{\lambda}$.
Figure 6: False color and contour plots of the stream function in the viscous regime for $\beta=3$, $l_{G}/R_{\mathrm{no}}=10$. (a) $\Psi(x,y)$ computed using Eqs. (89) and (97). (b) Enlarged view of the no-go region, computed using Eqs. (97) and (99).
In Fig. 6 we show the 2D plots of the stream function at intermediate distances $a\ll r\lesssim l_{G}$. The zoomed-in view of the no-go region [Fig. 6(b)] reveals spiral-like pattern of the current flow. The spiraling direction on the way from $y=+\infty$ to $y=$-$\infty$ is opposite to that in Fig. 3 because $\sin\theta_{L}<0$ in the viscous case whereas $\sin\theta_{L}>0$ in the diffusive case.
@@ IV Discussion
The exponentially small current inside the no-go region exhibits an intriguing spiral pattern. The spiraling is driven by the magnetic Lorentz force, so it has a different origin from vortical flow that arises in viscous hydrodynamics [72, 73, 74, 45, 75, 56] in the absence of the magnetic field. The twisting of the current lines at short distances has an effect on the asymptotic behavior. It determines the angular orientation of the Landauer dipole, i.e., the correction to the uniform current at large distances.
The Landauer dipole and the effective obstacle size we have investigated are the simplest and the most robust characteristics. They can be extracted already from STP maps collected at a moderate resolution. By varying $B$ and tracking how the no-go radius grows, one can in principle distinguish diffusive and hydrodynamic behavior and estimate the electron viscosity. High-resolution scans [22] have been shown to also resolve a small anomaly of the electrochemical potential predicted to occur [76, 77] at the distance of one cyclotron diameter $2R_{c}$ from the hard-wall barrier. In this regard, our classical continuum medium theory has a number of limitations. Both the diffusive and the hydrodynamic approximations become inaccurate at short distances. They do not account for features on the scale of the cyclotron radius where the motion is ballistic. These approaches also fail at low temperatures and long distances, where quantum localization — the hallmark of the QHE — becomes important. To remedy these shortcomings, self-consistent solutions of the Boltzmann kinetic equation and quantum transport simulations in realistic potential profiles may be necessary. The scaling rules we have derived may provide a useful reference for these more ambitious calculations.
@@ Acknowledgements
We are grateful to Z. Krebs and V. Brar for prior collaboration on Ref. [22] that inspired this study. We thank D. G. Polyakov for illuminating comments on the manuscript.
@@ Appendix A Exact solutions for the diffusive transport
$\lambda=\Lambda(\alpha,\beta)a^{2},$ (77)
$\Lambda(\alpha,\beta)=-\frac{\Gamma\left(1-\frac{2}{\beta}\right)}{\Gamma\left(1+\frac{2}{\beta}\right)}\,\frac{\Gamma\left(-\frac{i}{\alpha\beta}+\frac{1}{\beta}\right)}{\Gamma\left(-\frac{i}{\alpha\beta}-\frac{1}{\beta}\right)}\,.$ (78)
@@
The exact solution of Eq. (20) can be also obtained for the density profile
$n(r)=n_{0}\exp\left[\left({a}/{r}\right)^{\beta}\right],$ (79)
which models a density accumulation (Sec. II.4). Although $n(r)$ goes to infinity at the origin, it does not matter because this divergence is located deep inside the no-go region. Indeed, a physically acceptable solution for $\phi(r)$ is given by Eq. (76) with $a^{\beta}$ replaced by $a^{\beta}$. This potential and the corresponding electric current are exponentially small inside the no-go region and are regular at $r=0$ if Eq. (77) is replaced by
$\lambda\to\frac{1-i\alpha}{1+i\alpha}\,\lambda^{*}=\frac{1-i\alpha}{1+i\alpha}\,\Lambda(-\alpha,\beta)a^{2},$ (80)
where the star denotes the complex conjugation. If the magnetic field is strong enough, $1/\alpha\gg\beta$, then $\theta_{L}\simeq-\pi+\pi/\beta$ for the density accumulation case, as mentioned in Sec. II.4. To illustrate its dependence on the magnetic field, we plot $\theta_{L}$ as a function of $1/\alpha$ in Fig. 7 for both density depletion and accumulation cases.
Figure 7: Landauer dipole orientation angle vs. $1/\alpha$ for $\beta=3$. The red solid line is for the density depletion, Eq. (13). The black solid line is for the density accumulation, Eq. (79). The dashed lines show the asymptotic values $\pm\pi(1-1/\beta)$.